OpenStudy (anonymous):

Is the line through points P(0, –9) and Q(2, –8) perpendicular to the line through points R(1, 4) and S(3, 3)? Explain. Yes; their slopes are equal. Yes; their slopes have product –1. No; their slopes are not reciprocals. Yes; their slopes have product –1.

OpenStudy (anonymous):

The formula for slope is \[slope = (y_2-y_1)/(x_2-x_1)\] so the slope of PQ is:\[(-8-(-9))/(2-0) = 1/2\] The slope of RS is : \[(3-4)/(3-1) = -1/2\] In order for two slopes to be perpendicular, their product must be -1 (1/2)(-1/2) = -1/4 which isnt -1, so the lines arent perpendicular. Answer is C